# Stability Analysis of Grünwald Interpolation Operators on Chebyshev Nodes

### v1

@@ 4 Quantitative estimates
In this section, we obtain a quantitative estimate for the convergence of the sequence of operators $\{\mathcal{G}_{n}\}$ using both the first and second moduli of smoothness. We employ $K$-functionals as intermediate tools to derive these estimates. Such quantitative convergence results have been extensively studied for positive linear operators; see, for example, [4, 5, 18].
The classical modulus of smoothness of first and second orders and definition of K-functionals are given below. We also recall the equivalence of the K-functionals and the modulus of smoothness (see [4, 5]).
@@
and the second-order modulus of smoothness of $f$ by
$\omega_{2}(f,\delta)=\sup\{|f(x+2h)-2f(x+h)+f(x)|:x\in[a,b-2h],\;0<h\leq\delta\}.$
For $i=1,2$, the corresponding $K$-functional is defined by
$K_{i}(f,\delta)=\inf\{\big(\|f-g\|_{\infty}+\delta^{i}\|g^{(i)}\|_{\infty}:g\in C^{i}[0,1]\big),\quad\delta>0\},$
where $\|.\|_{\infty}$ denotes the sup-norm on $[0,1]$.
For $f\in C[a,b]$, there exist constants $C_{1},C_{2}>0$, depending only on $i$, such that the following equivalence holds (see [5]):
$C_{1}\,\omega_{i}(f,\delta)\leq K_{i}(f,\delta)\leq C_{2}\omega_{i}(f,\delta),\quad\delta>0.$
In the following theorem $\|.\|_{\infty}$ denotes the sup-norm on $0,\pi]$.
@@
$\|\mathcal{G}_{n}(f)-f\|_{\infty}\leq\frac{c_{2}+1}{2}\{C_{1}\omega(f,\mu_{n})+C_{2}\omega_{2}(f,\nu_{n})\}.$
where $C_{1}$ and $C_{2}$ are constants which does not depend on $n$ or $f$, $c_{2}$ is as in Lemma 2.4 and
$\mu_{n}=(\frac{\pi}{2n}+\frac{1}{n^{\frac{1}{3}}})\frac{c_{2}}{c_{2}+1}+\frac{9\pi^{4}}{64(c_{2}+1)}\frac{1}{n^{\frac{2}{3}}},$
$\nu_{n}=\frac{1}{2}\{(\frac{\pi^{2}}{4n^{2}}+\frac{1}{n^{\frac{2}{3}}}+\frac{\pi}{n^{\frac{4}{3}}})\frac{c_{2}}{c_{2}+1}+\frac{9\pi^{5}}{64(c_{2}+1)}\frac{1}{n^{\frac{2}{3}}}\}.$
@@
Let $g\in C^{2}[0,\pi]$ and $t,\theta\in[0,\pi]$. Then we have
$g(t)-g(\theta)=g^{\prime}(\theta)(t-\theta)+\int\limits_{\theta}^{t}g^{\prime\prime}(u)(t-u)\ du.$
We have
$|\int\limits_{\theta}^{t}g^{\prime\prime}(u)(t-u)\ du|\leq\|g^{\prime\prime}\|_{\infty}\int\limits_{\theta}^{t}|t-u|\ du\leq\|g^{\prime\prime}\|_{\infty}\frac{(t-\theta)^{2}}{2}.$
By the linearity of $\mathcal{G}_{n}$,
$\begin{split}\mathcal{G}_{n}(g)(\theta)-g(\theta)=g^{\prime}(\theta)\mathcal{G}_{n}(t-\theta)+\mathcal{G}_{n}(\int\limits_{\theta}^{t}g^{\prime\prime}(u)(t-u)\ du)(\theta).\end{split}$
Set $\epsilon_{n}=\frac{1}{n^{\frac{1}{3}}}$. Then,
$|\mathcal{G}_{n}(t-\theta)(\theta)|\leq\frac{1}{2}\sum\limits_{k=1}^{n}|\tilde{\theta}_{k}^{(n)}-\theta||l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\leq(\frac{\pi}{2n}+\epsilon_{n})c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}},$
and
$\begin{split}|\mathcal{G}_{n}(\int\limits_{\theta}^{t}g^{\prime\prime}(u)(t-u)\ du)(\theta)|&\leq\frac{1}{2}\sum\limits_{k=1}^{n}|\int\limits_{\theta}^{t}g^{\prime\prime}(u)(\tilde{\theta}_{k}^{(n)}-u)\ du)||l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\\ &\leq\frac{\|g^{\prime\prime}\|_{\infty}}{4}\sum\limits_{k=1}^{n}(\tilde{\theta}_{k}^{(n)}-\theta)^{2}|l_{k}(\theta-\frac{\pi}{2n})+l_{k}(\theta+\frac{\pi}{2n})|\\ &\leq\frac{\|g^{\prime\prime}\|_{\infty}}{2}\{(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}\}.\end{split}$
Thus,
$\|\mathcal{G}_{n}(g)-g\|_{\infty}\leq\|g^{\prime}\|_{\infty}\{(\frac{\pi}{2n}+\epsilon_{n})c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}}\}+\frac{\|g^{\prime\prime}\|_{\infty}}{2}\{(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}\}.$
Suppose $f\in C[0,\pi]$. Then for all $g\in C^{2}[0,1]$,
$\mathcal{G}_{n}(f)-f=\mathcal{G}_{n}(f-g)+\mathcal{G}_{n}(g)-g-(f-g).$
Taking sup-norm,
$\begin{split}\|\mathcal{G}_{n}(f)-f\|_{\infty}&\leq\|\mathcal{G}_{n}(f-g)\|_{\infty}+\|\mathcal{G}_{n}(g)-g\|_{\infty}+\|f-g\|_{\infty}\\ &\leq(c_{2}+1)\|f-g\|_{\infty}+\|\mathcal{G}_{n}(g)-g\|_{\infty}\\ &\leq(c_{2}+1)\|f-g\|_{\infty}+\|g^{\prime}\|_{\infty}\{(\frac{\pi}{2n}+\epsilon_{n})c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}}\}\\ &\quad\quad\quad+\frac{\|g^{\prime\prime}\|_{\infty}}{2}\{(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})c_{2}+\frac{9\pi^{5}}{64}\frac{1}{n\epsilon_{n}^{2}}\}\\ &=\frac{c_{2}+1}{2}\{\|f-g\|_{\infty}+\|g^{\prime}\|_{\infty}\{(\frac{\pi}{2n}+\epsilon_{n})\frac{c_{2}}{c_{2}+1}+\frac{9\pi^{4}}{64(c_{2}+1)}\frac{1}{n\epsilon_{n}^{2}}\}+\\ &\quad\quad\quad\|f-g\|_{\infty}+\frac{\|g^{\prime\prime}\|_{\infty}}{2}\{(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})\frac{c_{2}}{c_{2}+1}+\frac{9\pi^{5}}{64(c_{2}+1)}\frac{1}{n\epsilon_{n}^{2}}\}\}.\end{split}$
Let $\mu_{n}=(\frac{\pi}{2n}+\epsilon_{n})\frac{c_{2}}{c_{2}+1}+\frac{9\pi^{4}}{64(c_{2}+1)}\frac{1}{n\epsilon_{n}^{2}}$ and $\nu_{n}=\frac{1}{2}\{(\frac{\pi^{2}}{4n^{2}}+\epsilon_{n}^{2}+\frac{\pi\epsilon_{n}}{n})\frac{c_{2}}{c_{2}+1}+\frac{9\pi^{5}}{64(c_{2}+1)}\frac{1}{n\epsilon_{n}^{2}}\}$.
Hence we have,
$\|\mathcal{G}_{n}(f)-f\|_{\infty}\leq\frac{c_{2}+1}{2}\{\|f-g\|_{\infty}+\|g^{\prime}\|_{\infty}\mu_{n}+\|f-g\|_{\infty}+\|g^{\prime\prime}\|_{\infty}\nu_{n}\}.$
Now taking infimum over all $g\in C^{2}[0,1]$, we have,
$\|\mathcal{G}_{n}(f)-f\|_{\infty}\leq\frac{c_{2}+1}{2}\{K_{1}(f,\mu_{n})+K_{2}(f,\nu_{n})\}.$
By the equivalence of K-functionals and the modulus of continuity, we have
$\|\mathcal{G}_{n}(f)-f\|_{\infty}\leq\frac{c_{2}+1}{2}\{C_{1}\omega(f,\mu_{n})+C_{2}\omega_{2}(f,\nu_{n})\},$
for some positive constants $C_{1}$ and $C_{2}$. ∎

### v2

@@ 4 Quantitative estimates
In this section, we obtain a quantitative estimate for the convergence of the sequence of operators $\{\mathcal{G}_{n}\}$ using modulus of continuity. Such quantitative convergence results have been extensively studied for positive linear operators; see, for example, [4, 5, 18].
We recall the definition of the classical modulus of continuity.
@@
In the following, we obtain a quantitative estimate for $\{\mathcal{G}_{n}\}$ using modulus of continuity. Let $\|\cdot\|_{\infty}$ denotes the supremum norm on $0,\pi]$.
@@
$\|\mathcal{G}_{n}(f)-f\|_{\infty}\leq(c_{2}+1)\omega(f,\mu_{n}),$
where
$\mu_{n}:=\left(\frac{\pi}{2n}+\frac{1}{n^{\frac{1}{3}}}\right)c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n^{\frac{1}{3}}}.$
@@
Let $f\in C[0,\pi]$, $t,\theta\in[0,\pi]$ and $\delta>0$, then by the properties of the modulus of continuity,
$\begin{split}|f(t)-f(\theta)|&\leq\omega(f,|t-\theta|)\\ &\leq(1+|t-\theta|\delta^{-1})\omega(f,\delta)\end{split}$
Now,
$\begin{split}\mathcal{G}_{n}(f)(\theta)-f(\theta)&=\frac{1}{2}\sum\limits_{k=1}^{n}(f(\theta_{k}^{(n)})-f(\theta))\Big[l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\Big].\end{split}$
Thus
$|\mathcal{G}_{n}(f)(\theta)-f(\theta)|$ $\leq\frac{1}{2}\sum_{k=1}^{n}\bigl|f(\theta_{k}^{(n)})-f(\theta)\bigr|\bigl|\,l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\bigr|$
$\leq\frac{1}{2}\sum_{k=1}^{n}\Bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\Bigr|\,\omega(f,\delta)$
$\quad+\frac{1}{2}\sum_{k=1}^{n}\bigl|\theta_{k}^{(n)}-\theta\bigr|\Bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\Bigr|\delta^{-1}\,\omega(f,\delta)$
$\leq c_{2}\,\omega(f,\delta)+\frac{1}{2}\sum_{k=1}^{n}\bigl|\theta_{k}^{(n)}-\theta\bigr|\Bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\Bigr|\delta^{-1}\,\omega(f,\delta).$
Let
$\delta\;=\delta_{n}=\frac{1}{2}\sum_{k=1}^{n}\bigl|\theta_{k}^{(n)}-\theta\bigr|\Bigl|l_{k}(\theta-\tfrac{\pi}{2n})+l_{k}(\theta+\tfrac{\pi}{2n})\Bigr|\leq\left(\frac{\pi}{2n}+\epsilon_{n}\right)c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}}.$
The last inequality has been derived in Theorem 3.2, where $\epsilon_{n}=\frac{1}{n^{\frac{1}{3}}}$. Let
$\mu_{n}:=\left(\frac{\pi}{2n}+\epsilon_{n}\right)c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n\epsilon_{n}^{2}}=\left(\frac{\pi}{2n}+\frac{1}{n^{\frac{1}{3}}}\right)c_{2}+\frac{9\pi^{4}}{64}\frac{1}{n^{\frac{1}{3}}}.$
Hence, $|\mathcal{G}_{n}(f)(\theta)-f(\theta)|\leq(c_{2}+1)\omega(f,\mu_{n})$ for all $\theta\in[0,\pi]$. The result follows.